OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? A.3, 4, 5 B.28, 26, 12 C.17, 14, 6 D.all of the above

OpenStudy (texaschic101):

use that pythagorean theorem...A^2 + B^2 = C^2 just make sure, when subbing in your numbers, that C is the biggest number Now lets find our right triangles...I will do the first one for you.. A. (3,4,5)...A = 3, B = 4, C = 5 A^2 + B^2 = C^2 3^2 + 4^2 = 5^2 9 + 16 = 25 25 = 25 (correct).....these equal, therefore, they are the sides of a right triangle. Now see what you can do for the others

OpenStudy (anonymous):

can someone pleasssse help!

OpenStudy (anonymous):

so is it D?

Mehek (mehek14):

do you know how to use the pythagorean theorem? @chantel32 @texaschic101 literally explained what you have to do

OpenStudy (anonymous):

yes i do but i got it so thanks